OpenStudy (anonymous):

find the range of the given function f(x)=(x-[x])/(1+x-[x])

OpenStudy (anonymous):

is that floor function?

OpenStudy (anonymous):

Dont know

OpenStudy (anonymous):

[x] ? i think its greatest integer smaller than or equal to x

OpenStudy (anonymous):

do u know properties of that function?

OpenStudy (anonymous):

It is greatest integer function

OpenStudy (anonymous):

like [3.5]=3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

[-2.1] ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

answer to my question

OpenStudy (anonymous):

so u must know that \(0\le x-[x] < 1\)

OpenStudy (anonymous):

I am a mathematician so i know about it

mathslover (mathslover):

@Joseph91 if some one is helping you so be nice to him or leave hopes from us to be polite with you ...

mathslover (mathslover):

You are lucky that mukushla is helping you .. he is great in mathematics but he can only explain you if you will be nice to him

OpenStudy (anonymous):

hint: let \(t=x-[x]\) then u have \[f(t)=\frac{t}{1+t}\] with \(0 \le t \le1\) i hope u can find range of later function

OpenStudy (anonymous):

sorry \(0 \le t <1\)

OpenStudy (anonymous):

so what you had done nothing on it

OpenStudy (anonymous):

\[f(t)=1-\frac{1}{1+t}\]\[ 1 \le 1+t<2\]now what?

OpenStudy (anonymous):

[0,?)

OpenStudy (anonymous):

it is not the answer sir the answer is [0,0.5)

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

so see this \[0 \le t<1\]\[1 \le t+1<2\]\[\frac{1}{2} < \frac{1}{1+t}\le1\]\[-1 \le -\frac{1}{1+t}<-\frac{1}{2}\]\[0 \le 1-\frac{1}{1+t}<\frac{1}{2}\]\[0 \le f(t)<\frac{1}{2}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

cheking

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

so who are you sir? you solved it very easily!

OpenStudy (anonymous):

sorry about the questions... i just wanted to know what is your level